On average, four students visits the Mathematics and Statistics tutoring centre during a 5-minute period.

(a) Calculate the probability that three students visit the centre during a 5 minute period. (3)

(b) During a ten-minute period, what is the probability that at least 4 students visit the centre? (5)


1
Expert's answer
2022-04-27T00:44:18-0400

Let "X" be the random variable, "X =" number of students visiting the centre during t-minute period

"X\sim Po(\lambda t)"

(a)

"\lambda t=4"


"P(X=3)=\dfrac{e^{-4}\cdot(4)^3}{3!}=0.195367"

(b)

"\lambda t=8"


"P(X\ge4)=1-P(X<4)"

"=1-\dfrac{e^{-8}\cdot(8)^0}{0!}-\dfrac{e^{-8}\cdot(8)^1}{1!}"

"-\dfrac{e^{-8}\cdot(8)^2}{2!}-\dfrac{e^{-8}\cdot(8)^3}{3!}"

"=0.957620"